OpenStudy (anonymous):

help me please find an equation of the line passing through the point (6,5) and the prependicular to the line whose equation is 2y+3x=6

OpenStudy (amistre64):

switch your numbers and negate one

OpenStudy (amistre64):

\[2y+3x \rightarrow 3y -2x\]

OpenStudy (amistre64):

now plug in your point

OpenStudy (amistre64):

3(5)-2(6) = ? 15 -12 = 3 3y -2x = 3

OpenStudy (anonymous):

thats it?

OpenStudy (amistre64):

that it